OpenStudy (anonymous):

The area of a rectangle is x2 - 2x - 15 and the length of the rectangle is x + 3. Find the width of the rectangle.

OpenStudy (anonymous):

since Area = L*W then L = A/W so, x^2-2x-15 = (x+3)*W and W = (x^2-2x-15)/(x+3) you'll need to factor the area expression to (x-5)(x+3) now simplify...